### v2

@@ Abstract
In this paper, we present Export3D, a one-shot 3D-aware portrait animation method that is able to control the facial expression and camera view of a given portrait image. To achieve this, we introduce a tri-plane generator that directly generates a tri-plane of 3D prior by transferring the expression parameter of 3DMM into the source image. The tri-plane is then decoded into the image of different view through a differentiable volume rendering. Existing portrait animation methods heavily rely on image warping to transfer the expression in the motion space, challenging on disentanglement of appearance and expression. In contrast, we propose a contrastive pre-training framework for appearance-free expression parameter, eliminating undesirable appearance swap when transferring a cross-identity expression. Extensive experiments show that our pre-training framework can learn the appearance-free expression representation hidden in 3DMM, and our model can generate 3D-aware expression controllable portrait image without appearance swap in the cross-identity manner.
@@ 1 Introduction
In this paper, we address the appearance-expression entanglement issue by proposing a contrastive pre-training framework over video datasets that produces appearance-free facial expressions with orthogonal structure. Armed with this representation, we build a one-shot 3D-aware portrait image animation method, namely Export3D, which controls the facial expression and 3D camera view of a given source image without appearance swap. To achieve this, we design a generator architecture consisting of vision transformer (ViT) and convolution layers [56, 17, 44] that directly generates the tri-planes from the source image and driving expression parameters instead of predicting the deformation fields. The main contributions of this work are summarized as follows:
@@
  • We propose a contrastive pre-training framework for the appearance-free facial expression latents distilled from the 3DMM parameters where they form an orthogonal structure for different facial expressions.
@@ 2.2 Portrait Image Animation
Portrait image animation, or face reenactment, is a task that animates a given source image according to the input driving condition, either audio [45, 31, 41, 69, 37, 22] or image [52, 70, 57, 58, 64]. Specifically, image-driven methods transfer the motion of the driving image into the source image by learning the motion between them. Most works [52, 70, 57] use facial key-points as a pivot representation to be aware of motion via the key-point displacement. FOMM [52] estimates facial key-points in an unsupervised manner, approximating the motion through the first-order Taylor expansion. LIA [58] encodes a motion in terms of latent codes by introducing an orthonormal basis as a motion dictionary. However, the local motion (e.g., facial expression) and the global motion (e.g. head pose) are still entangled in those representations. DPE [42] proposes a bidirectional cyclic training strategy to decouple the pose and expression within the latent codes, while it produces video-level artifacts due to the instability of the cycle-consistency learning.
@@
Figure 1: Training overview of Export3D. We convert a source image $S\in\mathbb{R}^{3\times H\times W}$ into a tri-plane $T_{\beta_{D}}(S)$ for rich 3D priors, conditioned on an expression parameter $\beta_{D}\in\mathbb{R}^{64}$ from a driving image $D$. A differentiable volume rendering renders the tri-plane into a raw rendered image $\hat{D}_{raw}$ using the camera parameter $p_{D}\in\mathbb{R}^{25}$ of $D$, which is then super-resolved into a final image $\hat{D}$
@@ 3 Methods
First of all, we formulate our portrait animation method, Export3D. Given a source image $S\in\mathbb{R}^{3\times H\times W}$, our method transfers the facial expression and camera view of a driving image $D\in\mathbb{R}^{3\times H\times W}$ with the expression and camera parameters, respectively. Our method uses a tri-plane [10] as the intermediate feature representation, providing a 3D position information for differentiable volume rendering [40, 36]. We directly generate an expression-transferred tri-plane from the source image and the driving expression parameter [8, 3] through expression adaptive layer normalization (EAdaLN) (Sec. 3.2). Based on the observation that the expression parameter still contains the appearance information, we propose a pre-training framework using contrastive learning to obtain the appearance-free expression latents, which forms an orthogonal structure for different expressions (Sec. 3.1). The expression-transferred tri-plane is rendered into a 3D-aware image through the differentiable volume rendering, and then super-resolved into the final output (Sec. 3.3).
@@ 3.1 Contrastive Learned Basis Scaling (CLeBS)
where $f_{I}(\cdot)$ is an image encoder, $f_{e}(\cdot)$ is an expression encoder, $\tau$ is the temperature, and $\cos(\cdot,\cdot)$ is the cosine similarity, respectively. Since all samples are from the same video, they share the same appearance, thereby the objective (1) enforces the encoders to learn appearance-free expression.
Moreover, for designing the expression encoder $f_{e}(\cdot)$, we focus on the orthogonal structure of 3DMM [8] that controls different expressions along different orthogonal directions. To provide the appearance-free expression with the orthogonal structure [58], we introduce a learned orthonormal basis $V$:
@@
spanning our new expression sub-space ($\delta_{ij}$ is the Kroneker delta function). More precisely, we convert the expression $\beta\in\mathbb{R}^{64}$ into the low-dimensional coefficient $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\in\mathbb{R}^{n}~{}(n\ll 64)$ and then scales the learned orthonormal basis $V\in\mathbb{R}^{d}$ to produce the appearance-free expression representation $\beta^{\prime}\in\mathbb{R}^{d}$:
@@
In this space, an expression is a linear combination of the basis $V=\{\textbf{v}_{i}\}_{i=1}^{n}$ where the coefficient $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ is responsible for the intensity of each expression direction. We call our encoder $f_{e}(\cdot)$ a learned basis scaling (LeBS) module with contrastive pre-training (CLeBS). Note that once CLeBS is pre-trained with the objective (1), no further training is required as illustrated in Fig. 1 and Fig. 3.
@@ 3.2 Hybrid Tri-plane Generator
where $\text{T}_{ij}\in\mathbb{R}^{32\times\frac{H}{2}\times\frac{W}{2}}$ is the 32-dimensional feature of $\frac{H}{2}\times\frac{W}{2}$ resolution on the $ij$-plane. EG3D [10] utilizes StyleGAN2 [28] to generate the tri-plane from a noise, forming the style latent space $\mathcal{W}\subseteq\mathbb{R}^{512}$. Several works [62, 65, 54, 38, 33, 7, 66] extend the 2D GAN-inversion methods [1, 2, 48, 55, 49] to 3D GAN-inversion in terms of reconstructing the tri-plane from style latent code. However, these methods often face challenges in preserving facial identity since the style latent code lacks the capacity for encoding spatial information and person-specific visual details.
We directly generate the expression-transferred tri-plane $\text{T}_{\beta_{D}}(S)$ from the source $S$ and the driving expression $\beta_{D}\in\mathbb{R}^{64}$ to reconstruct the driving $D$. Inspired by Live3DPortrait [56], we construct the tri-plane generator with ViT and convolution [17, 61]. Specifically, we convert $S\in\mathbb{R}^{3\times H\times W}$ into a visual feature in $\mathbb{R}^{\frac{h}{4}\times\frac{H}{2^{3}}\times\frac{W}{2^{3}}}$ through a stack of convolutional blocks, and then merge it into the $h$-dimensional $\frac{H\cdot W}{2^{8}}$ visual tokens through a overlap patch merge operator [61]. These tokens and driving expression are processed through a conditional ViT [17, 44, 61] blocks, namely EAdaLN-ViT, where the expression modulates [44] the visual tokens through expression adaptive layer normalization (EAdaLN) as illustrated in Fig. 3. EAdaLN is applied right before the multi-head self-attention (MSA) and the mix feed-forward network (Mix-FFN)[61] of each ViT block to inject the semantic expression into the visual tokens:
@@
where $x$ is the input visual token, $\text{LN}(\cdot)$ is the layer normalization [5], $\sigma(\beta^{\prime}_{D})$ and $\mu(\beta^{\prime}_{D})$ are the $h$-dimensional scale and shift factors computed from $\beta^{\prime}_{D}=f_{e}(\beta_{D})\in\mathbb{R}^{d}$, respectively.
Figure 3: Hybrid tri-plane generator $\mathbf{G}$ and Expression Adaptive Layer Normalization (EAdaLN).
To efficiently propagate the visual tokens to the higher resolution, we up-sample the visual tokens with pixel shuffle [56] followed by the Gaussian low-pass filter [27]. We experimentally find that the tokens and the pixel shuffle produce grid artifacts, challenging to eliminate in the image space. Employing low-pass filters effectively mitigates these artifacts by smoothing the borderline artifacts over the coordinate. Lastly, we use ViT and convolutional blocks to output the tri-plane $\text{T}_{\beta_{D}}(S)$:
@@ 3.3 Volume Rendering and Super-resolution
The tri-plane can be rendered into a 2D RGB image through the differentiable volume rendering [40, 36, 10]. The expression-transferred tri-plane $\text{T}_{\beta}(S)$ is projected onto 3 orthogonal planes ($xy,yz,zx$-planes) and then aggregated through average [10]:
$F_{\beta}(S)=\frac{1}{3}(F_{\beta,xy}(S)+F_{\beta,yz}(S)+F_{\beta,zx}(S)),$ (7)
where $F_{\beta,ij}(S)$ are the projected features of $T_{\beta}(S)$ onto the $ij$ planes. A Light-weight MLP assigns a color $c$ and density $\mathbf{\sigma}$ to each point $(x,y,z)$ using the aggregated feature $F_{\beta}(S)$:
$\text{MLP}:F_{\beta}(S)\longrightarrow(c,\mathbf{\sigma}).$ (8)
@@
Note that the ray $\mathbf{r}$ is determined by the driving camera parameter $p_{D}\in\mathbb{R}^{25}$ to render the generated tri-plane $\text{T}_{\beta_{D}}(S)$ into a image of the same view with $D$. The appearance and expression are determined within the tri-plane generation process, thereby the volume rendering applies the pose the to produce 2D images of different views.
Directly rendering a target high-resolution image requires high computational cost. One promising approach to address this issue is to incorporate super-resolution blocks [10, 60, 56] that up-sample the rendered image of low resolution. Following this approach, we first render a $\hat{D}_{raw}\in\mathbb{R}^{3\times\frac{H}{4}\times\frac{W}{4}}$ and then apply the super-resolution to obtain the target resolution $\hat{D}\in\mathbb{R}^{3\times H\times W}$, as illustrated in Fig. 1. Instead of using style-modulated convolution [10, 56], we use plane convolutional blocks for super-resolution, as we do not leverage the style latent code. Detailed architecture is provided in supplementary materials.
3.4 Training Objectives
Our model is trained with reconstruction manner that reconstruct a driving frame $D$ from a source frame $S$ with the driving expression parameters $\beta_{D}$ and camera parameters $p_{D}$ where these frames are randomly sampled from the same video clip. The training consists of two stages. In the first phase, we employ MSE loss $\mathcal{L}_{2}$ and VGG16 [53] multi-scale perceptual loss $\mathcal{L}_{lpips}$ [68] to minimize the perceptual distance between the generated frame $\hat{D}$ and the driving frame $D$. We also minimize the distance between the raw rendered image $\hat{D}_{raw}$ and raw driving image $D_{raw}$ using the same loss functions, denoted by $\mathcal{L}^{raw}_{2}$ and $\mathcal{L}^{raw}_{lpips}$, respectively:
$\mathcal{L}_{rec}=\mathcal{L}^{raw}_{2}+\mathcal{L}_{2}+\mathcal{L}^{raw}_{lpips}+\mathcal{L}_{lpips}.$ (11)
In the second phase, we integrate the conditional discriminator used in [26], using the camera parameter as additional condition and employing binary cross-entropy loss to compute adversarial loss $\mathcal{L}_{adv}$. The total loss function $\mathcal{L}_{total}$ is
$\mathcal{L}_{total}=\lambda_{rec}\mathcal{L}_{rec}+\lambda_{adv}\mathcal{L}_{adv},$ (12)
where $\lambda_{rec}$ and $\lambda_{adv}$ are balancing coefficients.
@@ 4 Experiments
4.1 Dataset and Pre-processing
We train our model on real video dataset VFHQ [63]. Following the video pre-processing strategies in [52, 31], we convert the original video into 25 fps and crop the facial regions of resolution $256\times 256$, ensuring that the nose is located at the center of the frame. We use a 3DMM extractor [16] to obtain the expression parameter $\beta\in\mathbb{R}^{64}$. We adopt the pre-preprocesing strategy of EG3D [10] for the camera parameter $p\in\mathbb{R}^{25}$ (the concatenation of the camera intrinsic parameters in $\mathbb{R}^{9}$ and the inverse extrinsic parameter in $\mathbb{R}^{16}$). After the video pre-processing, 6196 video clips are used for training, and 50 videos are used for test. We also evaluate our model on the test dataset of TalkingHead-1KH [57]. After the same pre-processing, remaining 20 videos of different identities are used. Note that we do not use TalkingHead-1KH for training or fine-tuning.
4.2 Implementation Details
We set $n=10$ basis vectors for LeBS in the subspace in $\mathbb{R}^{256}$ ($d=256$). For contrastive pre-training of LeBS, we draw 32 negative samples per 1 positive sample and the temperature $\tau$ in (1) is $0.07$. We empirically set the balancing coefficients in (12) by $\lambda_{rec}=1$, and $\lambda_{adv}=0.01$. For all training, we use Adam [32] optimizer with the learning rate $10^{-4}$ for Export3D, $10^{-4}$ for CLeBS, and $10^{-5}$ for the discriminator, respectively. We train our model for 300,000 steps with the reconstruction loss (11) and then incorporate the adversarial loss (12) for 10,000 steps to slightly improve the visual quality. Overall training conducts on a single A100 GPU about 5 days with batch size 8. In the inference phase, we use randomly sampled frontal frame as the source frame.
4.3 Evaluation
Table 1: Quantitative comparison of on VFHQ. The best score for each metric is in bold. Note that we only measure CSIM [14], AED and APD [16, 47] for the cross-identity experiment as no ground-truth is available.
@@ 4.3 Evaluation
Table 2: Quantitative comparison of on TalkingHead-1KH. The best score for each metric is in bold. Note that we only measure CSIM [14], AED and APD [16, 47] for the cross-identity experiment as no ground-truth is available.

### v3

@@ Abstract
In this paper, we present Export3D, a one-shot 3D-aware portrait animation method that is able to control the facial expression and camera view of a given portrait image. To achieve this, we introduce a tri-plane generator with an effective expression conditioning method, which directly generates a tri-plane of 3D prior by transferring the expression parameter of 3DMM into the source image. The tri-plane is then decoded into the image of different view through a differentiable volume rendering. Existing portrait animation methods heavily rely on image warping to transfer the expression in the motion space, challenging on disentanglement of appearance and expression. In contrast, we propose a contrastive pre-training framework for appearance-free expression parameter, eliminating undesirable appearance swap when transferring a cross-identity expression. Extensive experiments show that our pre-training framework can learn the appearance-free expression representation hidden in 3DMM, and our model can generate 3D-aware expression controllable portrait images without appearance swap in the cross-identity manner.
@@ 1 Introduction
In this paper, we address the appearance-expression entanglement issue by proposing a contrastive pre-training framework over video datasets that produces appearance-free facial expressions with an orthogonal structure. Armed with this representation, we build a one-shot 3D-aware portrait image animation method, namely Export3D, which controls the facial expression and 3D camera view of a given source image without appearance swap. To achieve this, we design a generator architecture consisting of vision transformer (ViT) and convolution layers [56, 17, 44] that directly generates the tri-planes from the source image and driving expression parameters. Instead of predicting the deformation fields for the expression, we introduce an expression adaptive layer normalization (EAdaLN) which can effectively transfer the driving expression to the source image. The main contributions of this work are summarized as follows:
@@
  • We propose a contrastive pre-training framework for the appearance-free facial expression distilled from the 3DMM parameters where they form an orthogonal structure for different facial expressions.
@@ 2.2 Portrait Image Animation
Portrait image animation, or face reenactment, is a task that animates a given source image according to the input driving condition, either audio [45, 31, 41, 70, 37, 22] or image [52, 71, 58, 59, 65]. Specifically, image-driven methods transfer the motion of the driving image into the source image by learning the motion between them. Most works [52, 71, 58] use facial key-points as a pivot representation to be aware of motion via the key-point displacement. FOMM [52] estimates facial key-points in an unsupervised manner, approximating the motion through the first-order Taylor expansion. LIA [59] encodes a motion in terms of latent codes by introducing an orthonormal basis as a motion dictionary. However, the local motion (e.g., facial expression) and the global motion (e.g., head motion) are still entangled in those representations. DPE [42] proposes a bidirectional cyclic training strategy to decouple the pose and expression within the latent codes, while it produces video-level artifacts due to the instability of the cycle-consistency learning.
@@
Figure 1: Training overview of Export3D. We convert a source image $S\in\mathbb{R}^{3\times H\times W}$ into a tri-plane $T_{\beta_{D}}(S)$ for rich 3D priors, conditioned on an expression parameter $\beta_{D}\in\mathbb{R}^{64}$ from a driving image $D\in\mathbb{R}^{3\times H\times W}$. A differentiable volume rendering renders the tri-plane into a raw rendered image $\hat{D}_{raw}\in\mathbb{R}^{3\times\frac{H}{4}\times\frac{W}{4}}$ using the camera parameter $p_{D}\in\mathbb{R}^{25}$ of $D$, which is then super-resolved into a final image $\hat{D}\in\mathbb{R}^{3\times H\times W}$.
@@ 3 Methods
First of all, we formulate our portrait animation method, Export3D. Given a source image $S\in\mathbb{R}^{3\times H\times W}$, our method transfers the facial expression and camera view of a driving image $D\in\mathbb{R}^{3\times H\times W}$ with the expression and camera parameters, respectively. We employ a tri-plane [10] as the intermediate feature representation, providing a strong 3D position information for differentiable volume rendering [40, 36]. We directly generate an expression-transferred tri-plane from the source image and the driving expression parameter [8, 3] through expression adaptive layer normalization (EAdaLN) (Sec. 3.2). Based on the observation that the expression parameter still contains the appearance information, we propose a pre-training framework using contrastive learning to obtain the appearance-free expression, which forms an orthogonal structure for different expressions (Sec. 3.1). The expression-transferred tri-plane is rendered into a 3D-aware image through the differentiable volume rendering, and then super-resolved into the final output (Sec. 3.3).
@@ 3.1 Contrastive Learned Basis Scaling (CLeBS)
where $f_{I}(\cdot)$ is an image encoder, $f_{e}(\cdot)$ is an expression encoder, $\tau$ is the temperature, and $\cos(\cdot,\cdot)$ is the cosine similarity, respectively. Since all samples are from the same video, they share the same appearance, thereby Eq. 1 enforces the encoders to learn appearance-free expression.
Moreover, for designing the expression encoder $f_{e}(\cdot)$, we focus on the orthogonal structure of 3DMM [8] that controls different expressions along different orthogonal directions. To provide the appearance-free expression with the orthogonal structure, we introduce a learned orthonormal basis $V$:
@@
spanning our new expression sub-space ($\delta_{ij}$ is the Kroneker delta function). More precisely, we convert the expression $\beta\in\mathbb{R}^{64}$ into the low-dimensional coefficient $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\in\mathbb{R}^{n}~{}(n\ll 64)$ and then scales the learned orthonormal basis $V\subseteq\mathbb{R}^{d}$ to produce the appearance-free expression representation $\beta^{\prime}\in\mathbb{R}^{d}$:
@@
We apply QR-decomposition [59] to a learned weight ($\in\mathbb{R}^{d\times n}$) to explicitly compute the orthonormal basis $V\in\mathbb{R}^{d\times n}$. In this space, an expression is a linear combination of the basis $V=\{\textbf{v}_{i}\}_{i=1}^{n}$ where the coefficient $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ is responsible for the intensity of each expression direction. We call our encoder $f_{e}(\cdot)$ a learned basis scaling (LeBS) module with contrastive pre-training (CLeBS). Once CLeBS is pre-trained with Eq. 1, no further training is required as illustrated in Fig. 1 and Fig. 3, and the image encoder $f_{I}(\cdot)$ is never used after then.
@@ 3.2 Hybrid Tri-plane Generator
where $\text{T}_{ij}\in\mathbb{R}^{32\times\frac{H}{2}\times\frac{W}{2}}$ is the 32-dimensional feature of $\frac{H}{2}\times\frac{W}{2}$ resolution on the $ij$-plane. EG3D [10] utilizes StyleGAN2 [28] to generate the tri-plane from a noise, forming the style latent space $\mathcal{W}\subseteq\mathbb{R}^{512}$. Several works [63, 66, 54, 38, 33, 7, 67] extend the 2D GAN-inversion methods [1, 2, 48, 55, 49] to 3D GAN-inversion in terms of reconstructing the tri-plane from the style latent code.
Figure 3: Hybrid tri-plane generator $\mathbf{G}$ and Expression Adaptive Layer Normalization (EAdaLN). EAdaLN modulates the expression of $S$ using the refined expression $\beta^{\prime}$ from CLeBS.
However, these methods often face challenges in preserving facial identity since the style latent code lacks the capacity for encoding spatial information and person-specific visual details. We directly generate the expression-transferred tri-plane $\text{T}_{\beta_{D}}(S)$ from the source $S$ and the driving expression $\beta_{D}\in\mathbb{R}^{64}$ to reconstruct the driving $D$. Inspired by Live3DPortrait [56], we construct the tri-plane generator with ViT and convolution [17, 62]. Specifically, we convert $S\in\mathbb{R}^{3\times H\times W}$ into a visual feature in $\mathbb{R}^{\frac{h}{4}\times\frac{H}{2^{3}}\times\frac{W}{2^{3}}}$ through a stack of convolutional blocks, and then merge it into the $h$-dimensional $\frac{H\cdot W}{2^{8}}$ visual tokens through a overlap patch merge operator [62]. These tokens and driving expression are processed through a conditional ViT [17, 44, 62] blocks, namely EAdaLN-ViT, where the expression modulates [44] the visual tokens through expression adaptive layer normalization (EAdaLN) as illustrated in Fig. 3. EAdaLN is applied right before the multi-head self-attention (MSA) and the mix feed-forward network (Mix-FFN)[62] of each ViT block to inject the semantic expression into the visual tokens:
@@
where $x$ is the input visual token, $\text{LN}(\cdot)$ is the layer normalization [5], $\sigma(\beta^{\prime}_{D})$ and $\mu(\beta^{\prime}_{D})$ are the $h$-dimensional scale and shift factors computed from $\beta^{\prime}_{D}=f_{e}(\beta_{D})\in\mathbb{R}^{d}$, respectively. To efficiently propagate the visual tokens to the higher resolution, we upsample the visual tokens with pixel shuffle [56] followed by the Gaussian low-pass filter [27]. We experimentally find that the tokens and the pixel shuffle produce grid artifacts, challenging to eliminate in the image space. Employing low-pass filters effectively mitigates these artifacts by smoothing the borderline artifacts over the coordinate. Lastly, we use ViT and convolutional blocks to output the tri-plane $\text{T}_{\beta_{D}}(S)$:
@@ 3.3 Volume Rendering and Super-resolution
The tri-plane can be rendered into a 2D RGB image through the differentiable volume rendering [40, 36, 10]. The expression-transferred tri-plane $\text{T}_{\beta_{D}}(S)$ is projected onto 3 orthogonal planes ($xy,yz,zx$-planes) and then aggregated through average [10]:
$F_{\beta_{D}}(S)=\frac{1}{3}(F_{\beta_{D},xy}(S)+F_{\beta_{D},yz}(S)+F_{\beta_{D},zx}(S)),$ (7)
where $F_{\beta_{D},ij}(S)$ are the projected features of $T_{\beta_{D}}(S)$ onto the $ij$ planes. A light-weight MLP assigns a color $c$ and density $\mathbf{\sigma}$ to each point $(x,y,z)$ using the aggregated feature $F_{\beta_{D}}(S)$:
$\text{MLP}:F_{\beta_{D}}(S)\longrightarrow(c,\mathbf{\sigma}).$ (8)
@@
Note that the ray $\mathbf{r}$ is determined by the driving camera parameter $p_{D}\in\mathbb{R}^{25}$ to render the generated tri-plane $\text{T}_{\beta_{D}}(S)$ into a image of the same view with $D$. As the appearance and the expression are already encoded in the tri-plane generation, the volume rendering can determine the view-consistent images.
Directly rendering a target high-resolution image requires high computational cost. One promising approach to address this issue is to incorporate super-resolution blocks [10, 61, 56] that upsamples the rendered image of low resolution. Following this approach, we first render a $\hat{D}_{raw}\in\mathbb{R}^{3\times\frac{H}{4}\times\frac{W}{4}}$ and then apply the super-resolution to obtain the target resolution $\hat{D}\in\mathbb{R}^{3\times H\times W}$, as illustrated in Fig. 1. Instead of using style-modulated convolution [10, 56], we use plane convolutional blocks for super-resolution, as we do not leverage the style latent code. Detailed architecture is provided in supplementary materials.
@@ 4 Experiments
Table 1: Quantitative comparison on VFHQ. The best score for each metric is in bold. Note that we only measure CSIM [14], AED and APD [16, 47] for the cross-identity experiment as no ground-truth is available.
@@
4.1 Dataset and Pre-processing
We train our model on real video dataset VFHQ [64]. Following the video pre-processing strategies in [52, 31], we convert the original video into 25 fps and crop the facial regions of resolution $256\times 256$, ensuring that the nose is located at the center of the image. We use a 3DMM extractor [16] to obtain the expression parameter $\beta\in\mathbb{R}^{64}$. We adopt the pre-preprocesing strategy of EG3D [10] for the camera parameter $p\in\mathbb{R}^{25}$ (the concatenation of the camera intrinsic parameters in $\mathbb{R}^{9}$ and the inverse extrinsic parameter in $\mathbb{R}^{16}$). After the video pre-processing, 6196 video clips are used for training, and 50 videos are used for test. We also evaluate our model on the test dataset of TalkingHead-1KH [58]. After the same pre-processing, remaining 20 videos of different identities are used.
4.2 Evaluation
Table 2: Quantitative comparison on TalkingHead-1KH. The best score for each metric is in bold. Note that we only measure CSIM [14], AED and APD [16, 47] for the cross-identity experiment as no ground-truth is available.